OpenStudy (anonymous):

If 4.0 × 10^-18 J of work done on a point charge changes its electric potential difference by 25.0 V, then what is the magnitude of the charge? A. 1.6 × 10^-19 C B. 4.0 × 10^-19 C C. 1.6 × 10^-18 C D. 4.0 × 10^-18 C

OpenStudy (michele_laino):

the requested charge (magnitude) is: \[q = \frac{W}{{\Delta V}}\]

OpenStudy (anonymous):

ok! 4.0*10^-18 / 25 ?

OpenStudy (michele_laino):

\[q = \frac{W}{{\Delta V}} = \frac{{4 \times {{10}^{ - 18}}}}{{25}}\]

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

yay! so we get 1.6E-19? so choice A is our solution?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

yay!! thanks!!:D

OpenStudy (michele_laino):

:)